# LIR: Efficient Degradation Removal for Lightweight Image Restoration

### 02 Feb 2024 v1

@@ Abstract
Recently, there have been significant advancements in Image Restoration based on CNN and transformer. However, the inherent characteristics of the Image Restoration task are often overlooked in many works. These works often focus on the basic block design and stack numerous basic blocks to the model, leading to redundant parameters and unnecessary computations and hindering the efficiency of the image restoration. In this paper, we propose a Lightweight Image Restoration network called LIR to efficiently remove degradation (blur, rain, noise, haze, etc.). A key component in LIR is the Efficient Adaptive Attention (EAA) Block, which is mainly composed of Adaptive Filters and Attention Blocks. It is capable of adaptively sharpening contours, removing degradation, and capturing global information in various image restoration scenes in an efficient and computation-friendly manner. In addition, through a simple structural design, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks. Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art networks on most benchmarks with fewer parameters and computations. It is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic. The code is available at: https://github.com/Dongqi-Fan/LIR
@@ 1 Introduction
Image restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, denoising, dehazing, etc. In recent years, we have witnessed the great process of deep learning methods [30, 61, 53, 35, 41, 55, 71] in IR. However, many existing works focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task. This approach often leads to parameter redundancy and unnecessary computations. Instead, our goal is an efficient and lightweight network design for various IR tasks.
@@
To address the issue of vanishing gradient and stabilize training, as shown in Fig. 1 (a), most modern networks usually incorporate local and global residual connections. However, we think that these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output. To validate our idea, we conduct visualization experiments on the deblurring and deraining tasks using Restormer [62], a transformer model with a U-Net structure. Fig. 1 illustrates the feature maps output by the encoder and decoder at three levels, as well as the input and output images of the Restormer, on the deblurring (at the top of Fig. 1 (d)) and deraining task (at the bottom of Fig. 1 (d)). Notably, at the bottom of Fig. 1 (d), it is evident that each feature map and the final output of the network (Fig. 1 (b)) all contain degradation (rain). Although less pronounced in the deblurring task, a similar phenomenon can be observed at the top of Fig. 1 (d). To overcome this issue, we use an ingenious structural design (Fig. 2) that effectively eliminates the degradation in the local and global residual connections. (Fig. 1 (c)) shows the restored image output by our LIR, which is more plain than the Restormer (Fig. 1 (b)).
@@
Moreover, we observe another serious issue in the visualization. In the visualization of the deblurring task (at the top of Fig. 1 (d)), the feature maps generated by encoder_level2 and decoder_level2 exhibit clear contours while the other feature maps exhibit blurred contours. This issue is even worse in the deraining task (at the bottom of Fig. 1 (d)). All feature maps not only display very blurry contours but also contain a considerable amount of noise. Especially for the feature map generated by decoder_level1, it shows barely discernible contours, only with faint traces of rain. The more important is that the high-frequency information has been lost. These observations indicate that although many modern networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and fails to address the core issues of image restoration, thus resulting in poor visual quality. Based on this insight, we propose the Adaptive Filter (Fig. 3) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
@@
Transformer-based models have demonstrated impressive performance in various tasks, both at high-level [12, 9, 33, 49] and low-level [6, 3, 55, 52, 29]. However, there exist different opinions. [11] argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs). In addition, [31, 48, 50] share a similar viewpoint that highlights the potential of MLPs beyond transformers. On the other hand, [51] suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy. Thus, excessive computations can be avoided. Furthermore, [16, 34, 8, 73] take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer. Satisfactory results are achieved and excessive computations are avoided. Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module that aims to capture global information of the features and keeps parameters and computations friendly.
@@
In this paper, we propose a Lightweight Image Restoration network called LIR (Fig. 2) to efficiently remove degradation and adapt to various IR tasks. The key component of LIR is the novel Efficient Adaptive Attention (EAA) Block consisting of several Adaptive Filters and Attention Blocks. Thanks to the Adaptive Filter and Attention Block, the EAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information. Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and EAA Block. This subtraction allows us to obtain the degradation, which is then expanded through the Transpose operation. After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection. Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art networks in PSNR and produces better visual results that are more in line with the human aesthetic.
@@
  • We introduce the EAA Block, which is not only able to adaptively enhance the object contour and remove degradation but also capture global information.
  • Building upon the EAA Block, we propose the LIR, a method that efficiently removes degradation and adapts to different IR tasks, achieving comparable performance to state-of-the-art networks and better visual effects with smaller parameters and computations.
Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Efficient Adaptive Attention (EAA) Blocks. Thanks to the Adaptive Filter, Patch Attention, and Channel Attention modules, the EAA Block is endowed with the ability to extract a variety of information efficiently.
@@ 2.2 Vision Transformer for Image Restoration
IPT [3] is the first work that introduces ViT [12] to low-level tasks to achieve image super-resolution, denoising, and deraining. The computational complexity of the self-attention part in ViT is $O(N^{2}D)$, where $N=H$$\times$$W$ represents the length of the patch sequence and $D$ represents the dimension of the sequence. Although vision transformers are capable of capturing long-range dependencies, the self-attention component imposes a significant computing burden. The computational complexity of the self-attention increases quadratically with the input size. To mitigate this issue, TransWeather [54] introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$. Uformer [55] and StripFormer [52] proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size. To further reduce the computational complexity, [23] estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL [29] proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention. The computational complexities are reduced to $O(NlogND)$ and $O(NMD)$ respectively, where M represents the size of feature map. However, the computation required for self-attention remains unaffordable. To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings in practical applications.
@@ 3.1 Lightweight Image Restoration
The overall structure of LIR is illustrated in Fig. 2, which primarily consists of Efficient Adaptive Attention (EAA) Blocks. Both LIR and EAA Block adopt a down-up style to reduce computation and save parameters, which downsizing the feature map in the beginning and upsizing at the end. To preserve important low-frequency information and ensure stable training, we add two global residual connections into LIR. One connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv. As mentioned in Sec. 1, these residual connections can bring degradation. To address this, we subtract the output of the fourth EAA Block from the output of the first EAA Block to obtain the corresponding degradation. We then amplify the degradation through the Transpose operation and subtract it from the residual connection to obtain a clean connection. The same approach is used to remove degradation in the local residual connection within the EAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation. The reason we adopt this method is that the feature map of the shallow layer contains more degradation compared to a deep layer. Thus, this design allows the network to dynamically perceive this difference between shallow and deep layers during the training process and inference stage. Next, we illustrate the Adaptive Filter and Attention Block in detail.
@@ 3.2 Adaptive Filter
To enhance the object contour, remove degradation, and extract useful high-frequency information, inspired by Edge-oriented Convolution Block (ECB) [69] (at the right of Fig. 3), we develop the Adaptive Filter (AF) (at the left of Fig. 3). The AF shares similarities with the ECB, but significantly outperforms it in three ways. Firstly, the ECB introduces more convolution operations during training leading to more training costs. Instead, our AF only requires 10 training parameters resulting in smaller training costs. Secondly, the AF can be applied to various IR scenes and dynamically adjust the dependence on filters by adjusting parameters $\alpha$. However, the ECB treats each branch equally and can not adapt to various IR scenes. Lastly, the AF offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
@@
Each filter in the AF has its own advantages and disadvantages. We use the learnable parameter $\alpha$ to maximize their advantages and minimize their disadvantages in different IR scenarios. Specifically, the Roberts filter is used to detect the pixel changes in the features by the differential method. It is effective in extracting prominent edges but fails to detect details. Where Roberts-x and Roberts-y represent the detection in the horizontal direction and vertical direction; The Sobel filter extracts edge information by approximating first-order differentials discretely. It has some resistance to noise and a good ability to detect details. However, the Sobel filter may cause edge thinning and false detection in scenes with a mass of noise; The Laplacian filter extracts edge information by approximating second-order differentials discretely. It can detect more detailed edges and textures than Sobel but is more sensitive to noise and may result in a bilateral effect on the edges. Where Laplacian-4 and Laplacian-8 represent 4 domains and 8 domains; The high-pass filter obtains high-frequency information by subtracting the features after Gaussian blur processing from the original features. This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the AF for stages where the features have scarce noise and a clear outline.
3.3 Attention Block
Our Attention Block consists of four components: ResBlock, Patch Attention, SimpleGate, and ResCABlock. We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module. The efficacy of the gate function has been demonstrated in prior research [31, 53, 5]. Here, we simply employ the SimpleGate module from [5] to further extract dependencies between features. Finally, we leverage the ResCABlock to extract channel-wise information.
ResBlock and ResCABlock. Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk. The only difference is that the latter has a channel attention module at the tail. The Channel Attention module [70], based on Squeeze-and-Excitation (SE) [17], has gained much popularity for its ability to extract channel-wise information. Many studies [29, 6, 5, 72] have directly utilized this module in their models. Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information.
Patch Attention. In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead. Instead, we simplify the attention map computation in our PA (Fig. 4) by using just one MLP and one element-wise multiplication. Specifically, we set C, H, and W as the channel, height, and width of the feature map, and $(P_{h},P_{w})$ as patch size. The $C_{scale}$ is used to control the dimension of the channel. We start by reducing the dimension of the input feature map from $(C,H,W)$ to $(C//C_{scale},H,W)$ using a $1\times 1$ convolution. Next, we divide the feature map into non-overlapping patches and reshape it into $(HW//P_{h}P_{w},P_{h}P_{w}C//C_{scale})$. Then, we feed this reshaped feature into the MLP to obtain the attention weights for each patch $(HW//P_{h}P_{w},1,1)$. Finally, we perform element-wise multiplication between the obtained attention weights and the input feature map to capture global information. It is important to note that our attention weights have a smaller dimension, resulting in a lower computation scale to self-attention. The $(P_{h},P_{w})$ in the LIR is set to $(8,8)$, $(4,4)$, and $(2,2)$ respectively, and the $C_{scale}$ is set to 16 and 8, according to the depth of the network.

### 29 Feb 2024 v2

# LIR: A Lightweight Baseline for Image Restoration

@@ Abstract
Recently, there have been significant advancements in Image Restoration based on CNN and transformer. However, the inherent characteristics of the Image Restoration task are often overlooked. Many works, instead, only focus on the basic block design and stack numerous such blocks to the model, leading to parameters redundant and computations unnecessary. Thus, the efficiency of the image restoration is hindered. In this paper, we propose a Lightweight Baseline for Image Restoration called LIR to efficiently reconstruct the image and remove degradations (blur, rain, noise, haze). First of all, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks, through a simple structural design. Then, to achieve lightweight, a Lightweight Adaptive Attention (LAA) Block is introduced depending on the inherent characteristics of the Image Restoration, which is mainly composed of proposed Adaptive Filters and Attention Blocks. LAA is capable of adaptively sharpening contours, removing degradation, and capturing global information in various Image Restoration scenes in a computation-friendly manner. Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art models with fewer parameters and computations in certain tasks. In addition, it is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic. The code is available at: https://github.com/Dongqi-Fan/LIR
Keywords:
Image Restoration Lightweight Attention
[…]
[width=1]figures/introduction.eps
@@ 1 Introduction
Image Restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, etc. In recent years, we have witnessed the great process of deep learning methods [28, 59, 52, 33, 39, 54, 67] in IR. However, many existing works only focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task. This approach often leads to parameter redundancy and computations unnecessary. Instead, our goal is an efficient and lightweight network design for various IR tasks.
@@
To address the issue of vanishing gradient and stabilize training, as shown in Fig. 1 (a), most modern networks usually incorporate local and global residual connections. However, we think that these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output. To validate our idea, we conduct visualization experiments on the deblurring and deraining tasks using Restormer [58], a transformer model with a U-Net structure. Fig. 1 illustrates the feature maps output by the encoder and decoder at three levels, as well as the input and output images of the Restormer, on the deblurring (at the top of Fig. 1 (d)) and deraining (at the bottom of Fig. 1 (d)) task. Notably, at the bottom of Fig. 1 (d), it is evident that each feature map and the final output of the network (Fig. 1 (b)) all contain degradation (rain). Although less pronounced in the deblurring task, a similar phenomenon can be observed at the top of Fig. 1 (d). To overcome this issue, we use an ingenious structural design (transpose branchs in the LIR and LAA Fig. 2) that effectively eliminates the degradation in the local and global residual connections. Fig. 1 (c) shows the restored image output by our LIR, which is more plain than the Restormer (Fig. 1 (b)).
@@
Moreover, we observe another serious issue in the visualization. In the visualization of the deblurring task (at the top of Fig. 1 (d)), the feature maps generated by encoder_level2 and decoder_level2 exhibit clear contours while the other feature maps exhibit blurred contours. This issue is even worse in the deraining (at the bottom of Fig. 1 (d)) task. All feature maps not only display very blurry contours but also contain a considerable amount of noise. Especially for the feature map generated by decoder_level1, it shows barely discernible contours, only with faint traces of rain. The more important is that the high-frequency information has been lost. These observations indicate that although many modern networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and fails to address the core issues of image restoration, thus resulting in poor visual quality. Based on this insight, we propose the Adaptive Filter (Fig. 3) (an component of the LAA) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
@@
Transformer-based models have demonstrated impressive performance in various tasks, both at high-level [12, 9, 31, 49] and low-level [6, 3, 54, 51, 27]. However, there exist different opinions. [10] argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs). In addition, [29, 47, 48] share a similar viewpoint that highlights the potential of MLPs beyond transformers. On the other hand, [50] suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy. Thus, excessive computations can be avoided. Furthermore, [16, 32, 8, 70] take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer. Satisfactory results are achieved and excessive computations are avoided. Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module (Fig. 4) (an component of the Attention Block) that aims to capture global information of the features and keeps parameters and computations friendly.
@@
In this paper, we propose a Lightweight Image Restoration network called LIR (Fig. 2) to efficiently remove degradation and adapt to various IR tasks. The key component of LIR is the novel Lightweight Adaptive Attention Block (LAA) consisting of several proposed Adaptive Filters and Attention Blocks. Futhermore, the Attention Block is formed by proposed Patch Attention module (Fig. 4), ResCABlock, and ResBlock (Details can be found in Fig. 2). Thanks to the Adaptive Filter and Attention Block, the LAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information. Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and LAA Block. This subtraction allows us to obtain the degradation, which is then expanded through the Transpose operation. After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection. Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art models in PSNR and produces better visual results that are more in line with the human aesthetic.
@@
  • Given the problem of the loss of high-frequency information observed in the visualization experiments, we propose the Adaptive Filter (Fig. 3) as a solution to adaptively extract high-frequency information and enhance object contours.
  • To give LIR the ability to extract global information, we propose the Patch Attention module (Fig. 4), which is more computation-friendly than the self-attention of the transformer.
  • Based on Adaptive Filter, Patch Attention, ResCABlock, and ResBlock, we designed a lightweight and powerful block: LAA. It can make the LIR achieve comparable performance to state-of-the-art networks and better visual effects with smaller parameters and computations.
@@ 2.2 Vision Transformer for Image Restoration
IPT [3] is the first work that introduces ViT [12] to low-level tasks to achieve image super-resolution, denoising, and deraining. The computational complexity of the self-attention part in ViT is $O(N^{2}D)$, where $N=H$$\times$$W$ represents the length of the patch sequence and $D$ represents the dimension of the sequence. Although vision transformers are capable of capturing long-range dependencies, the self-attention component imposes a significant computing burden. The computational complexity of the self-attention increases quadratically with the input size. To mitigate this issue, TransWeather [53] introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$. Uformer [54] and StripFormer [51] proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size. To further reduce the computational complexity, [22] estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL [27] proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention. The computational complexities are reduced to $O(NlogND)$ and $O(NMD)$ respectively, where M represents the size of feature map. However, the computation required for self-attention remains unaffordable. To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings.
[…]
[width=]figures/model_2.eps
Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Lightweight Adaptive Attention (LAA) Blocks. Thanks to the Adaptive Filter, Patch Attention, ResCABlock, and the ResBlock, the LAA Block is endowed with the ability to extract a variety of information efficiently.
@@ 3.1 Lightweight Image Restoration
The overall structure of LIR is illustrated in Fig. 2, which primarily consists of Lightweight Adaptive Attention (LAA) Blocks. Both LIR and LAA Block adopt a down-up style to reduce computation and save parameters, that is downsizing the feature map in the beginning and upsizing at the end. To preserve low-frequency information and ensure training is stable, we add two global residual connections into LIR. One connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv.
As mentioned in Sec. 1, these residual connections can bring degradation. To address this, we subtract the output of the fourth LAA Block from the output of the first LAA Block to obtain the corresponding degradation. We then amplify the degradation through the Transpose operation and subtract it from the residual connection to obtain a clean connection. The same approach is used to remove degradation in the local residual connection within the LAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation. The reason we adopt this method is that the feature map of the shallow layer contains more degradation compared to a deep layer. Thus, this design allows the network to dynamically perceive this difference between shallow and deep layers during the training process and inference stage. Now, the problem that local and global residual connections bring degradation is addressed. Next, we illustrate how we can achieve lightweight and efficiency through the Adaptive Filter and Attention Block in detail.
[…]
[width=]figures/filters_2.eps
@@ 3.2 Adaptive Filter
To achieve lightweight and efficiency, we first introduce the Adaptive Filter (at the left of Fig. 3), inspired by Edge-oriented Convolution Block (ECB) [66] (at the right of Fig. 3). The rich filters of different kinds allow the AF to enhance the object contour, remove degradation, and extract useful high-frequency information. Furthermore, the LAA block can be reparameterized as one $3\times 3$ and one high-pass filter after training to achieve lightweight.
Although Adaptive Filter shares similarities with the ECB, it significantly outperforms ECB in three ways. Firstly, the ECB introduces more convolution operations during training leading to more training costs (i.e. parameters and computations). Instead, our Adaptive Filter only requires 10 training parameters resulting in smaller training costs. Secondly, the Adaptive Filter can be applied to various IR scenes and dynamically adjust the dependence on different filters by adjusting parameters $\alpha$. However, the ECB treats each branch equally and can not adapt to various IR scenes. Lastly, the Adaptive Filter offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
Each filter in the Adaptive Filter has its own advantages and disadvantages. Thus, we use the learnable parameters $\alpha$ to adaptively maximize their advantages and minimize their disadvantages in different LAA blocks of the LIR. Specifically, in these filters, the Roberts filter is used to detect the pixel changes in the features by the differential method. It is effective in extracting prominent edges but fails to detect details. Where Roberts-x and Roberts-y represent the detection in the horizontal direction and vertical direction; The Sobel filter extracts edge information by approximating first-order differentials discretely. It has some resistance to noise and a good ability to detect details. However, the Sobel filter may cause edge thinning and false detection in scenes with a mass of noise; The Laplacian filter extracts edge information by approximating second-order differentials discretely. It can detect more detailed edges and textures than Sobel but is more sensitive to noise and may result in a bilateral effect on the edges. Where Laplacian-4 and Laplacian-8 represent 4 domains and 8 domains; The high-pass filter obtains high-frequency information by subtracting the features after Gaussian blur processing from the original features. This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the AF for stages where the features have scarce noise and a clear outline.
3.3 Attention Block
While different types of filters in Adaptive Filter (AF) enable LAA blocks to extract rich information, AF alone is not enough, as we need to take full advantage of the rich information extracted by AF. Thus, a strong and efficient Attention Block is introduced, which consists of four components: ResBlock, Patch Attention (PA), SimpleGate, and ResCABlock. We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module. The efficacy of the gate function has been demonstrated in prior research [29, 52, 4]. Here, we simply employ the SimpleGate module from [4] to further extract dependencies between features. Finally, we leverage the ResCABlock to extract channel-wise information.
3.3.1 ResBlock and ResCABlock
Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk. The only difference between ResBlock and ResCABlock is that the latter has a channel attention module at the tail. The Channel Attention module [68], based on Squeeze-and-Excitation (SE) [17], has gained much popularity for its ability to extract channel-wise information. Many studies [27, 6, 4, 69] have directly utilized this module in their models. Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information. In addition, the structure of ResBlock and ResCABlock are designed to be simple to meet lightweight and efficient requirements.
[…]
[width=]figures/patch_attention_2.eps
@@ 3.3.1 ResBlock and ResCABlock
3.3.2 Patch Attention
To give the Attention Block the ability to capture global information, we introduce the Patch Attention (PA) (Fig. 4). In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead. Instead, we simplify the attention map computation in our PA by using just one MLP and one element-wise multiplication.
Specifically, we set C, H, and W as the channel, height, and width of the feature map, and $(P_{h},P_{w})$ as patch size. The $C_{scale}$ is used to control the dimension of the channel. We start by reducing the dimension of the input feature map from $(C,H,W)$ to $(C//C_{scale},H,W)$ using a $1\times 1$ convolution. Next, we divide the feature map into non-overlapping patches and reshape it into $(HW//P_{h}P_{w},P_{h}P_{w}C//C_{scale})$. Then, we feed this reshaped feature into the MLP to obtain the attention weights for each patch $(HW//P_{h}P_{w},1,1)$. Finally, we perform element-wise multiplication between the obtained attention weights and the input feature map to capture global information. It is important to note that our attention weights have a smaller dimension, resulting in a lower computation scale to self-attention. The $(P_{h},P_{w})$ in the LIR is set to $(8,8)$, $(4,4)$, and $(2,2)$ respectively, and the $C_{scale}$ is set to 16 and 8, according to the depth of the network.
[…]
0.9 Method Type Rain100L [56] Params [M] GFLOPs PSNR $\downarrow$ SSIM UMRL [57] CNN 29.18 0.923 0.98 9.31 PReNet [42] RNN 32.44 0.950 0.17 37.27 AirNet [24] CNN 34.90 0.966 8.93 169.47 MPRNet [59] CNN 36.40 0.965 20.13 960.39 SPAIR [40] CNN 36.93 0.969 $\triangle$ $\triangle$ HINet [5] CNN 37.28 0.970 88.67 95.92 MAXIM-2S [52] MLP 38.06 0.977 22.50 80.13 SFNet [7] CNN 38.21 0.974 13.27 70.00 DRCNet [25] CNN 38.23 0.976 18.90 $\triangle$ Restormer [58] Transformer 38.99 0.978 26.13 79.31 LIR (Ours) CNN 38.96 0.983 7.31 63.96
[…]
[width=]figures/derain100l.eps